OpenStudy (anonymous):

Please help me with this I forgot how to do this :P ASAP!! solve the system { 2x +y - 4z = -30 { 3x + 2y +2z =-8 {5x + 5y + z = -26 A. (-30,-8,-26) B.(4,-2,4) C. (-4,-2,4) D.(-4,2,-4

OpenStudy (jack1):

{ 2x +y - 4z = -30 { 3x + 2y +2z =-8 {5x + 5y + z = -26 solve via matrix or simultaneous equations (ie substitution method)

OpenStudy (jack1):

{ 2x +y - 4z = -30 therefore y = -30 -2x +4z sub that into equation 2 { 3x + 2y +2z =-8 3x + 2(-30 -2x +4z) +2z =-8 3x + -60 -4x +8z +2z =-8 -x -60 + 10z =-8 x =-52 + 10z sub that into equation 3 {5x + 5y + z = -26 ...etc

OpenStudy (jack1):

does that make sense @rockstargirl17 ?

OpenStudy (anonymous):

yeah it kind of does

OpenStudy (jack1):

sweet, good luck